OpenStudy (anonymous):

Factor completely: 3x2 + 5x + 1 (3x + 1)(x + 1) (3x + 5)(x + 1) (3x − 5)(x + 1) Prime

OpenStudy (anonymous):

this expression is not factorable

OpenStudy (anonymous):

so it would be prime???

OpenStudy (anonymous):

yes! but more importantly, do you understand why?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

there are no 2 numbers that add up to 5 and multiply to 3

OpenStudy (anonymous):

is it because 3*3=9 so leaving you with 9x + 5x + 1 and 5 is your odd number in the equation which make the equation nonfactorable

OpenStudy (anonymous):

nope, thats not how you do it! when you factor in the form ax^2 + bx +c, you have to find two numbers that multiply to a*c and add up to b. So if those number dont exist, its not factorable. ex. 3x^2 + 5x + 1 a = 3, b =5, c = 1 a*c = 3 b = 5 what numbers multiply to 3? list the factors: 1,3 -- thats it 1 + 3 = 4. 4/=5! therefore, not factorable

OpenStudy (anonymous):

oh ok thank you for the help